OpenStudy (anonymous):

how do I write 4 logb^m + 1/2logb^n - 3logb^2p as one log?

OpenStudy (anonymous):

maybe if I draw this it will make more sense

OpenStudy (rajee_sam):

is the base to the log is b or is it b^m ( b to the power m)

OpenStudy (anonymous):

here let me just take a picture of the problem

OpenStudy (anonymous):

there

OpenStudy (anonymous):

Is that ok @rajee_sam ?

OpenStudy (rajee_sam):

Ok, hang on for the ride

OpenStudy (rajee_sam):

\[4\log_{b} m + \frac{ 1 }{ 2 }\log_{b} n - 3 \log_{b} 2p\]

OpenStudy (anonymous):

oh dear..

OpenStudy (rajee_sam):

do you know the basic log rules Log A + Log B = Log AB Log A - Log B = Log (A/B) m Log A = Log A^m

OpenStudy (rajee_sam):

all should have the same base

OpenStudy (rajee_sam):

here all logs have the same base which is 'b'

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm following you so far.

OpenStudy (rajee_sam):

now lets apply the rules one by one. first we need to simplify each log separately before we can combine them. For convenience I am not writing the b. since all bases are same I will write it at the end. But in your work you write it when you are solving in each and every step. Now, How can you rewrite 4 log m

OpenStudy (rajee_sam):

which rule can we apply?

OpenStudy (anonymous):

The first one?

OpenStudy (rajee_sam):

First one has two logs added. here we have only one ---> m Log A

OpenStudy (rajee_sam):

how can I rewrite this single log?

OpenStudy (anonymous):

ok, and you just moved the 4 to be the exponent of m?

OpenStudy (rajee_sam):

\[4 \log_{b} m = \log_{b} m ^{4}\]

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

yeah similarly how can I rewrite the other two independently

OpenStudy (rajee_sam):

\[\frac{ 1 }{ 2 }\log_{b} n = ?\]

OpenStudy (anonymous):

logb n^1/2 ?

OpenStudy (anonymous):

Sorry I don't know how to use the tools so well

OpenStudy (rajee_sam):

yes and the third one

OpenStudy (rajee_sam):

That's alright

OpenStudy (anonymous):

logb 2p^-3?

OpenStudy (rajee_sam):

\[\log_{b} m ^{4} + \log_{b} n ^{\frac{ 1 }{ 2 }} - \log_{b} (2p)^{3}\]

OpenStudy (rajee_sam):

I will keep the negative sign and just take the 3 to the power

OpenStudy (anonymous):

ok, fabulous. But I'm not done, right? I still need to write it as one log.

OpenStudy (rajee_sam):

not yet

OpenStudy (rajee_sam):

now we have to combine them two at a time

OpenStudy (rajee_sam):

once you are familiar with it you can do all at one go. But to understand better we will do two at a time now

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

Now to combine 1st and 2nd we have something like Log A + Log B , so using the rule how can write it as a single log?

OpenStudy (rajee_sam):

Log A + Log B = Log AB

OpenStudy (anonymous):

they are both log b, so that stays the same, right? And then do I just put m^4 and n^1/2 next to eachother?

OpenStudy (anonymous):

like logb m^4 n^1/2?

OpenStudy (rajee_sam):

yes. you are writing them together with a multiplication sign in between them Log A + Log B = Log AB ; AB actually means A x B

OpenStudy (rajee_sam):

Yes you are right

OpenStudy (rajee_sam):

\[\log_{b} m ^{4}n ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

YAY

OpenStudy (rajee_sam):

now again you have two logs this one you just wrote and the 3rd one which is of the form Log A - Log B = Log (A/B)

OpenStudy (rajee_sam):

now use the other rule to write them together

OpenStudy (anonymous):

ok so logb m^4 n^1/2 / logb (sp)^3

OpenStudy (anonymous):

I meant 2 not s

OpenStudy (rajee_sam):

no there should be only one log in your answer

OpenStudy (rajee_sam):

log A - log B = log (A/B)

OpenStudy (rajee_sam):

you have written log A / Log B

OpenStudy (rajee_sam):

\[\log_{b} \frac{ m ^{4} n ^{\frac{ 1 }{ 2 }} }{ (2p)^{3} }\]

OpenStudy (rajee_sam):

and voila that is your answer

OpenStudy (anonymous):

ahh ok awesome, thanks so much!!!!